# Kummer-Artin-Schreier-Witt Theory

### v2

@@ 6 The Group Structure
With the notations of Corollary 4.13, the special fiber of $\psi_{s}:\mathcal{W}_{s}(R)\rightarrow\mathcal{V}_{s}(R)$ is the Artin-Schreier-Witt isogeny
$W_{s}(\mathbb{F}_{p})\rightarrow W_{s}(\mathbb{F}_{p}),\quad(y_{0},y_{1}+b_{1},\ldots,y_{s-1}+b_{s-1})\mapsto\wp(y_{0},y_{1}+b_{1},\ldots,y_{s-1}+b_{s-1}).$
@@
We have a commutative diagram of group morphisms over $\hat{R}$:
[…]
Furthermore, removing the map $W_{s}\rightarrow W_{s}$ yields a commutative diagram of homomorphisms of group schemes over $R$.
@@
The commutativity of the first square follows directly from Definition 6.1, which defines the group structure via the Witt vector group structure.
For the second square, we define the maps $\tilde{E}_{s,p}$ and $\tilde{\Psi}_{s}$ as follows:
$\begin{split}(Y_{0}^{\prime},\ldots,Y_{s}^{\prime})&\mapsto(E_{1,p}((Y_{0}^{\prime})),\ldots,E_{s+1,p}((Y_{0}^{\prime},\ldots,Y_{s}^{\prime}))),\\ (X_{0}^{\prime},\ldots,X_{s}^{\prime})&\mapsto(G_{1,p}((X_{0}^{\prime})),\ldots,G_{s+1,p}((X_{0}^{\prime},\ldots,X_{s}^{\prime}))).\end{split}$
The fact that these maps are group homomorphisms follows from (13) and (14). Additionally, each $E_{j,p}((Y_{0}^{\prime},\ldots,Y_{j-1}^{\prime}))$ (resp. $G_{j,p}((X_{0}^{\prime},\ldots,X_{j-1}^{\prime}))$) is a polynomial in $\mathbb{Z}_{(p)}[\zeta_{p^{j}}][Y_{0},\ldots,Y_{j-1}]$ (resp. $\mathbb{Z}_{(p)}[\zeta_{p^{j}}][X_{0},\ldots,X_{j-1}]$). Consequently, the compositions $\tilde{E}_{s,p}\circ\gamma$ and $\tilde{\Psi}_{s}\circ\tau$ define morphisms of algebraic group schemes over $R$ that are isomorphic on the generic fibers.
Recall that $\phi_{s}$ is defined in (9). The commutativity of the second square follows from:
$\left(E_{i+1,p}((Y_{0}^{\prime},\ldots,Y_{i}^{\prime}))\right)^{p}\left(E_{i,p}((Y_{0}^{\prime},\ldots,Y_{i-1}^{\prime}))\right)^{-1}=Z_{i}^{p}Z_{i-1}^{-1}=G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i}^{\prime})).$
Finally, the commutativity of the third square has also been established in (9). This completes the proof of the commutativity of the entire diagram. ∎
@@ 7 The Universality
In this section, we establish the proof of Theorem 1.2. Our methodology follows standard procedures as outlined in [11, Section 3] and [7, Section 6.4]. To begin, we fix some notation. Let $R$ be a finite extension of $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$, and let $\pi$ be its uniformizer. Define $R_{\pi}:=R/\pi$, and denote by $\mathbb{G}_{m,R_{\pi}}$ the special fiber of $\mathbb{G}_{m}:=\operatorname{Spec}R\left[X,\frac{1}{X}\right]$. We have a natural closed immersion $\mathbb{G}_{m,R_{\pi}}\hookrightarrow\mathbb{G}_{m}$ denoted by $\iota$.
For the computation of $\mathcal{W}_{1}=\operatorname{Spec}R\left[Y_{0},\frac{1}{1+\lambda Y_{0}}\right]$, the following exact sequence of sheaves on the étale site of $\operatorname{Spec}R$ is particularly useful (see [10, Lemma 1.1]):
@@
where $\alpha:x\mapsto 1+\lambda x$ and $\beta:t\mapsto t\mod\pi$. The following result is derived from [11, Theorem 2.2].
@@
$0\to\mathcal{W}_{1}(B)\to\mathbb{G}_{m}(B)$ $=B^{*}\to\iota_{*}\mathbb{G}_{m,R_{\pi}}(B)=(B/(\pi B))^{*}\to$
$\to\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})\to\operatorname{H}^{1}(\operatorname{Spec}B,\mathbb{G}_{m})\to\cdots$
@@
00{0}${\mathbb{Z}_{p}/(p)}$${\mathbb{Z}_{p}/(p^{s})}$${\mathbb{Z}_{p}/(p^{s-1})}$00{0}00{0}${\mathcal{W}_{1}}$${\mathcal{W}_{s}}$${\mathcal{W}_{s-1}}$00{0}
@@
Moreover, the Artin-Schreier-Witt theory establishes the desired commutative diagram. ∎
@@
$0\rightarrow\mathcal{W}_{1}(B)\rightarrow\mathcal{W}_{s}(B)\rightarrow\mathcal{W}_{s-1}(B)\rightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})\rightarrow\\ \rightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s})\rightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s-1})\rightarrow\cdots$
@@
We will now proceed to prove our main theorem, which states that the group scheme $\mathcal{W}_{s}$ over $R$ provides a unified framework for the Kummer-Artin-Schreier-Witt theory.
@@
$0\rightarrow\mathbb{Z}/p^{s}\rightarrow\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}\rightarrow 0$
@@
$0\displaystyle 0\to\mathcal{V}_{s}(B)/\mathcal{W}_{s}(B)\rightarrow\mathrm{H}^{1}(\operatorname{Spec}B,\mathbb{Z}/p^{s})\rightarrow\mathrm{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s})=$ 0 .
@@
It implies that $\mathbb{Z}/p^{s}$-torsors over $\operatorname{Spec}B$ are in one-to-one correspondence with elements of $\mathcal{V}_{s}(B)/\mathcal{W}_{s}(B)$. ∎

### v3

@@ 6 The Group Structure
With the notations of Corollary 4.13, the special-fiber morphism $(\psi_{s})_{k}:\mathcal{W}_{s,k}\to\mathcal{V}_{s,k}$ is the Artin–Schreier–Witt isogeny
$W_{s}(\mathbb{F}_{p})\longrightarrow W_{s}(\mathbb{F}_{p}),\quad(y_{0},y_{1}+b_{1},\ldots,y_{s-1}+b_{s-1})\mapsto\wp(y_{0},y_{1}+b_{1},\ldots,y_{s-1}+b_{s-1}).$
@@
On formal completions along the identity sections, we have a commutative diagram of morphisms of formal group schemes over $R$:
$\widehat{\mathcal{W}}_{s}$$\widehat{W}_{s}$$^{\widehat{(\mathbb{G}_{m})^{s}}}$$\widehat{\mathcal{V}}_{s}$$\widehat{W}_{s}$$^{\widehat{(\mathbb{G}_{m})^{s}}}$$\scriptstyle{\widehat{\gamma}}$$\scriptstyle{\widehat{\psi}_{s}}$$\scriptstyle{\tilde{E}_{s,p}}$$\scriptstyle{\widehat{\phi}_{s}}$$\scriptstyle{\widehat{\tau}}$$\scriptstyle{\tilde{\Psi}_{s}}$
Here $\widehat{\phi}_{s}$ is the morphism on formal completions induced by $\phi_{s}$ from (10).
@@
The commutativity of the first square follows from the description of the restriction of $\psi_{s}$ to formal completions,
$\widehat{\psi}_{s}=\widehat{\tau}^{-1}\circ\widehat{\gamma}$,
(cf. Lemma 6.1), equivalently $\widehat{\tau}\circ\widehat{\psi}_{s}=\widehat{\gamma}$.
For the second square, we define the maps $\tilde{E}_{s,p}$ and $\tilde{\Psi}_{s}$ on $\widehat{W}_{s}$ by
$\begin{split}(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})&\mapsto(E_{1,p}((Y_{0}^{\prime})),\ldots,E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime}))),\\ (X_{0}^{\prime},\ldots,X_{s-1}^{\prime})&\mapsto(G_{1,p}((X_{0}^{\prime})),\ldots,G_{s,p}((X_{0}^{\prime},\ldots,X_{s-1}^{\prime}))).\end{split}$
The fact that these are homomorphisms of (formal) groups follows from (15) and (16).
Moreover, by construction of the formal coordinates (cf. (37) and (38)), we have
$Z_{i}=E_{i+1,p}((Y_{0}^{\prime},\ldots,Y_{i}^{\prime}))\qquad\text{and}\qquad U_{i}=G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i}^{\prime}))$
for $i=0,\ldots,s-1$.
Recall that $\phi_{s}$ is defined in (10). The commutativity of the second square follows from
$\left(E_{i+1,p}((X_{0}^{\prime},\ldots,X_{i}^{\prime}))\right)^{p}\left(E_{i,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime}))\right)^{-1}=G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i}^{\prime})).$
This is exactly the defining relation of $G_{i+1,p}$ in terms of $E_{i+1,p}$ (see the definition preceding (16)), together with $E_{i+1,p}(p\,\underline{a})=E_{i+1,p}(\underline{a})^{p}$ for Witt multiplication by $p$. This completes the proof. ∎
Proposition 6.13.
There is a commutative diagram of homomorphisms of algebraic group schemes over $R$:
$\mathcal{W}_{s}$$(\mathbb{G}_{m})^{s}$$\mathcal{V}_{s}$$(\mathbb{G}_{m})^{s}$$\scriptstyle{(Z_{0},\ldots,Z_{s-1})}$$\scriptstyle{\psi_{s}}$$\scriptstyle{\phi_{s}}$$\scriptstyle{(U_{0},\ldots,U_{s-1})}$
where $U_{0},\ldots,U_{s-1}\in\mathcal{O}(\mathcal{V}_{s})^{\times}$ and $Z_{0},\ldots,Z_{s-1}\in\mathcal{O}(\mathcal{W}_{s})^{\times}$ are the standard units from the definitions of $\mathcal{V}_{s}$ and $\mathcal{W}_{s}$. Here $\phi_{s}$ is as in (10). This diagram is the algebraic version of Proposition 6.12: one replaces the formal torus coordinates by the corresponding global units $(Z_{0},\ldots,Z_{s-1})$ and $(U_{0},\ldots,U_{s-1})$.
Proof.
The horizontal arrows are induced by the tuples of units $(Z_{0},\ldots,Z_{s-1})$ on $\mathcal{W}_{s}$ and $(U_{0},\ldots,U_{s-1})$ on $\mathcal{V}_{s}$. Moreover, by construction of the descended Hopf algebra structures on $\mathcal{O}(\mathcal{W}_{s})$ and $\mathcal{O}(\mathcal{V}_{s})$, each $Z_{i}$ and $U_{i}$ is group-like (i.e. $\Delta_{B}(Z_{i})=Z_{i}\otimes Z_{i}$ and $\Delta_{A}(U_{i})=U_{i}\otimes U_{i}$), hence these maps to $(\mathbb{G}_{m})^{s}$ are homomorphisms of group schemes.
For the left square, by construction of $\psi_{s}$ (cf. (35)), in $\mathcal{O}(\mathcal{W}_{s})^{\times}$ we have
$U_{0}=Z_{0}^{p},\qquad U_{j}=Z_{j}^{p}Z_{j-1}^{-1}\ \ (1\leq j\leq s-1).$
Equivalently, $(U_{0},\ldots,U_{s-1})=\phi_{s}(Z_{0},\ldots,Z_{s-1})$, so the square commutes. ∎
@@ 7 The Universality
In this section, we establish the proof of Theorem 1.2. Our methodology follows standard procedures as outlined in [21, Section 3] and [17, Section 6.4]. To begin, we fix some notation. Let $R$ be a finite extension of $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$, and let $\pi$ be its uniformizer. Define $R_{\pi}:=R/\pi$, and denote by $\mathbb{G}_{m,R_{\pi}}$ the special fiber of $\mathbb{G}_{m}:=\operatorname{Spec}R\left[X,\frac{1}{X}\right]$. We have a natural closed immersion $\mathbb{G}_{m,R_{\pi}}\hookrightarrow\mathbb{G}_{m}$ denoted by $\iota$.
For the computation of $\mathcal{W}_{1}=\operatorname{Spec}R\left[Y_{0},\frac{1}{1+\lambda Y_{0}}\right]$, the following exact sequence of sheaves on the étale site of $\operatorname{Spec}R$ is particularly useful (see [20, Lemma 1.1]):
@@
where $\alpha:x\mapsto 1+\lambda x$ and $\beta:t\mapsto t\mod\pi$. The following result is derived from [21, Theorem 2.2].
@@
$0\longrightarrow\mathcal{W}_{1}(B)\longrightarrow\mathbb{G}_{m}(B)$ $=B^{*}\longrightarrow\iota_{*}\mathbb{G}_{m,R_{\pi}}(B)=(B/(\pi B))^{*}\longrightarrow$
$\longrightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})\longrightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathbb{G}_{m})\longrightarrow\cdots$
@@
[…]
@@
Moreover, the Artin–Schreier–Witt theory establishes the desired commutative diagram. ∎
@@
$0\longrightarrow\mathcal{W}_{1}(B)\longrightarrow\mathcal{W}_{s}(B)\longrightarrow\mathcal{W}_{s-1}(B)\longrightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})\longrightarrow\\ \longrightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s})\longrightarrow\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s-1})\longrightarrow\cdots$
@@
We will now proceed to prove our main theorem, which states that the group scheme $\mathcal{W}_{s}$ over $R$ provides a unified framework for the Kummer–Artin–Schreier–Witt theory.
@@
$0\longrightarrow\mathbb{Z}/p^{s}\longrightarrow\mathcal{W}_{s}\longrightarrow\mathcal{V}_{s}\longrightarrow 0$
@@
$0\longrightarrow\mathcal{V}_{s}(B)/\mathcal{W}_{s}(B)\longrightarrow\mathrm{H}^{1}(\operatorname{Spec}B,\mathbb{Z}/p^{s})\longrightarrow\mathrm{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{s})=0.$
@@
It implies that $\mathbb{Z}/p^{s}$-torsors over $\operatorname{Spec}B$ are in one-to-one correspondence with elements of $\mathcal{V}_{s}(B)/\psi_{s}(\mathcal{W}_{s}(B))$. ∎
8 Matsuda’s Theory and the Refined Swan Conductor
Let $\mathbb{K}$ be a complete discretely valued field with residue field $\kappa$ of characteristic $p>0$ and uniformizer $\pi$. Let $e$ be the absolute ramification index of $\mathbb{K}$. When $\kappa$ is perfect, classical ramification theory studies the behavior of finite Galois extensions in great detail. One important invariant arising from this theory is the Swan conductor [22, Chapter IV], which measures the degree of wild ramification and has many applications in algebraic geometry.
For example, let $X$ be a curve over an algebraically closed field $\kappa$, let $U\subset X$ be a nonempty open subset, and let $\mathscr{F}$ be a $\mathbb{Q}_{\ell}$-sheaf on $U$. Then, for each closed point $x\in X\setminus U$, the Swan conductor $\mathrm{sw}_{x}(\mathscr{F})$ is a nonnegative integer measuring the wild ramification of the associated extension of $\widehat{\mathcal{O}}_{X,x}\cong\kappa\llbracket t\rrbracket$. The Grothendieck–Ogg–Shafarevich formula can be expressed in terms of $\mathrm{sw}_{x}(\mathscr{F})$ as follows:
$\chi_{c}(U_{\overline{\kappa}},\mathscr{F})=\operatorname{rank}(\mathscr{F})\cdot\chi_{c}(U_{\overline{\kappa}},\mathbb{Q}_{\ell})-\sum_{x\in X\setminus U}\mathrm{sw}_{x}(\mathscr{F}).$ (44)
When the residue field is not perfect, there is an analogous notion of the refined Swan conductor, which may be viewed as the Swan conductor together with additional differential data. Note that the Swan conductor vanishes precisely when there is no wild ramification; in particular, in our setting this implies that the associated residue field extension is separable. When the Swan conductor is positive, one can write
$\mathrm{rsw}(\mathscr{F})=\pi^{-e\,\mathrm{sw}(\mathscr{F})}\otimes\omega\in\mathfrak{p}_{\mathbb{K}}^{-\mathrm{sw}(\mathscr{F})}\otimes\Omega^{1}_{\kappa}.$
Recall that
$\mathfrak{p}_{\mathbb{K}}^{t}:=\{x\in\mathbb{K}\mid\nu(x)\geq t\}.$
The differential $\omega$ is usually called the differential Swan conductor of $\mathscr{F}$. Introduced by Kato in 1989 for characters of degree one, this invariant has since been thoroughly studied by various authors, including Deligne, Laumon, Saito, and Abbes. They have approached the subject from different perspectives and obtained numerous applications in a wide range of contexts (see [15], [11], [1], [24], [13], [2], [4]). For instance, formula (44) extends to higher dimensions [9] and gives rise to a vanishing cycle formula [7], a valuable tool for detecting good reduction.
The primary case of interest for the lifting problem is when $U$ is an open dense subset of the closed unit disc $\mathcal{D}=\operatorname{Spec}R\{X\}$, where $R$ is a complete discrete valuation ring with residue field $\kappa$. Let $\mathbb{K}$ be the function field of $\mathcal{D}$; its residue field $\kappa(x)$ is imperfect.
8.1 The equal-characteristic case
When $\mathbb{K}$ has equal characteristic $p>0$, the study of $\mathscr{F}\in\mathrm{H}^{1}(\mathbb{K},\mathbb{Q}/\mathbb{Z})$ often suffices to understand many properties of the Swan conductor in this setting. Recall that
$\mathrm{H}^{1}(\mathbb{K},\mathbb{Z}/p^{s})\cong W_{s}(\mathbb{K})\big/\wp\!\left(W_{s}(\mathbb{K})\right)$
(see §2.1). In this context, we have the refined Swan homomorphism [15]:
$\begin{split}\mathrm{Rsw}:\mathrm{Fil}_{n}W_{s}(\mathbb{K})&\longrightarrow\mathrm{Fil}_{n}\Omega^{1}_{\mathbb{K}}\big/\mathrm{Fil}_{\lfloor n/p\rfloor}\Omega^{1}_{\mathbb{K}},\\ (a_{0},\ldots,a_{s-1})&\longmapsto\omega:=\sum_{i=0}^{s-1}a_{i}^{p^{\,s-i-1}-1}\,da_{i},\end{split}$ (45)
where $n\in\mathbb{Q}_{\geq 0}$, and $\mathrm{Fil}$ on the left-hand side denotes Brylinski’s filtration of Witt vectors [5]. In particular, when $(a_{0},\ldots,a_{s-1})$ is best (see [12, Definition 2.2]), the refined Swan conductor is determined by $\omega$.
Theorem 8.1 ([8, Lemma 3.7]).
In the above notation, suppose $(a_{0},\ldots,a_{s-1})$ is best. Then
$\mathrm{rsw}(\mathscr{F})=\omega\otimes 1=\pi^{e\nu(\omega)}\otimes\overline{\pi^{-e\nu(\omega)}\sum_{i}a_{i}^{p^{\,s-i-1}-1}\,da_{i}}\in\mathfrak{p}_{\mathbb{K}}^{-\mathrm{sw}(\mathscr{F})}\otimes\Omega^{1}_{\kappa},$
where $\mathrm{sw}(\mathscr{F})=-\nu(\omega)$ is the Swan conductor.
This description makes it possible to compute and deduce numerous properties of $\mathrm{sw}(\mathscr{F})$; see [12], [26], [6], [3].
A length-one Witt vector $(a_{0})\in W_{1}(\mathbb{K})$ is best if $\nu(a_{0})<0$ and the reduction of $x_{0}:=\pi^{-e\nu(a_{0})}a_{0}$ is not a $p$-power.
Corollary 8.2.
In the above notation, suppose $\mathscr{F}\in\mathrm{H}^{1}(\mathbb{K},\mathbb{Z}/p)$ is represented by a best length-one Witt vector $(a_{0})\in\mathbb{K}$. Then
$\mathrm{rsw}(\mathscr{F})=da_{0}\otimes 1=\pi^{e\nu(a_{0})}\otimes d\overline{x}_{0}\in\mathfrak{p}_{\mathbb{K}}^{-\mathrm{sw}(\mathscr{F})}\otimes\Omega^{1}_{\kappa},$ (46)
where $\mathrm{sw}(\mathscr{F})=-\nu(a_{0})$.
Proof.
Apply Theorem 8.1 with $s=1$ and note that for a best length-one Witt vector we have $\omega=da_{0}$, $\nu(\omega)=\nu(a_{0})$, and $d\overline{x}_{0}=\overline{\pi^{-e\nu(a_{0})}da_{0}}$. ∎
8.2 The mixed-characteristic case
There is no known algorithm to compute the refined Swan conductor of a cyclic character in mixed characteristic $(0,p)$ using its Kummer class like Theorem 8.1. The only known case is for
$\mathscr{F}\in\mathrm{H}^{1}(\mathbb{K},\mathbb{Z}/p)\cong\mathbb{K}^{\times}/(\mathbb{K}^{\times})^{p}.$
Under certain mild assumptions, we may assume that the extension associated with $\mathscr{F}$ is fierce (see [25, §2.3]). We may then further assume that the corresponding Kummer element $u\in\mathbb{K}^{\times}$ is of one of the following forms [25, Proposition 5.2]:
  1. 1.
    $u\in\mathcal{O}_{\mathbb{K}}^{\times}$ and $\overline{u}\notin\kappa^{p}$.
  2. 2.
    $u=1+\alpha^{p}w$ with $\alpha\in\mathbb{K}$ satisfying $0<\nu(\alpha)<\tfrac{1}{p-1}$, and $w\in\mathcal{O}_{\mathbb{K}}$ such that $\overline{w}\notin\kappa^{p}$.
If $u$ is of one of the above forms, we say that it is reduced.
Proposition 8.3 ([25, Proposition 5.3]).
Suppose $\mathscr{F}=\mathfrak{K}_{p}(u)\in\mathrm{H}^{1}(\mathbb{K},\mathbb{Z}/p)$ where $u$ is reduced. Then the refined Swan conductor $\mathrm{rsw}(\mathscr{F})$ is given as follows.
  1. 1)
    If $\overline{u}\notin\kappa^{p}$ (Case 1), then
    $\mathrm{rsw}(\mathscr{F})=\lambda^{-p}\otimes\frac{d\overline{u}}{\overline{u}}.$
  2. 2)
    If $u=1+\alpha^{p}w$ (Case 2), then
    $\mathrm{rsw}(\mathscr{F})=\alpha^{p}\lambda^{-p}\otimes d\overline{w}.$
The following result recovers Corollary 8.2 in Case 2.
Proposition 8.4.
Suppose $\mathscr{F}=\mathfrak{K}_{p}(u)\in\mathrm{H}^{1}(\mathbb{K},\mathbb{Z}/p)$, where $u$ is reduced as in Case 2. Suppose moreover that $X_{0}\in\mathbb{K}$ is the Witt vector coordinate associated with $u$. Set $x_{0}:=X_{0}\,\pi^{-e\nu(X_{0})}$. Then
$\mathrm{rsw}(\mathscr{F})=-dX_{0}\otimes 1=\pi^{e\nu(X_{0})}\otimes\bigl(-d\overline{x}_{0}\bigr).$
Proof.
We are in the situation of Case 2 of reduced form, so $u=1+\alpha^{p}w$. The first Witt vector coordinate is given by
$X_{0}=\frac{1}{p\lambda}\ln(1+\alpha^{p}w).$
We expand the terms on the right-hand side to isolate the leading term. First, the expansion of the logarithm gives
$\ln(1+\alpha^{p}w)=\alpha^{p}w+h_{1},\qquad\nu(h_{1})\geq 2p\nu(\alpha).$
Second, consider the expansion $(1+\lambda)^{p}=1$, which gives $p\lambda+\lambda^{p}=-\sum_{k=2}^{p-1}\binom{p}{k}\lambda^{k}$. If $p=2$, then this identity simply reads $2\lambda+\lambda^{2}=0$, i.e. $p\lambda=-\lambda^{p}$. Assume now that $p>2$. The term with the lowest valuation on the right is $\binom{p}{2}\lambda^{2}$, which has valuation $1+\frac{2}{p-1}=\frac{p+1}{p-1}$. Thus, in all cases we have the relation
$p\lambda\equiv-\lambda^{p}\pmod{\mathfrak{p}_{\mathbb{K}}^{\frac{p+1}{p-1}}}$.
This implies
$\frac{1}{p\lambda}=-\lambda^{-p}+\delta,$ (47)
where the error $\delta$ has valuation strictly larger than $\nu(\lambda^{-p})$.
Multiplying these expansions, we find
$X_{0}$ $=(-\lambda^{-p}+\delta)(\alpha^{p}w+h_{1})$
$=-\lambda^{-p}\alpha^{p}w+\left(-\lambda^{-p}h_{1}\,+\,\delta\alpha^{p}w\,+\,\delta h_{1}\right).$
Let $C_{0}:=-\alpha^{p}\lambda^{-p}$. The main term $C_{0}w$ has valuation $p\nu(\alpha)-\frac{p}{p-1}$. Since $p\nu(\alpha)>0$, one checks easily that every term in the parenthesis has valuation strictly larger than $\nu(C_{0}w)$. Thus, we have the key valuation inequality
$\nu(X_{0}-C_{0}w)>\nu(C_{0}w)=\nu(X_{0}).$
Now define the normalized unit
$x_{0}:=X_{0}\pi^{-e\nu(X_{0})}.$
The valuation inequality above implies that the difference $X_{0}-C_{0}w$ has strictly larger valuation than $X_{0}$ itself. After multiplying by $\pi^{-e\nu(X_{0})}$, we obtain an element of positive valuation, hence it vanishes in the residue field. Therefore, in $\kappa$ we obtain the clean reduction
$\overline{x}_{0}=\overline{C}\,\overline{w},\qquad C:=C_{0}\,\pi^{-e\nu(X_{0})}\in\mathcal{O}_{\mathbb{K}}^{\times},$
where $C$ is a unit since $\overline{w}\neq 0$. Since $\overline{C}$ is a constant scalar in $\kappa^{\times}$, taking differentials gives the relation
$d\overline{x}_{0}=\overline{C}\,d\overline{w}.$
Substituting this into the expression for $\mathrm{rsw}(\mathscr{F})$ in the statement:
$\mathrm{rsw}(\mathscr{F})$ $=\pi^{e\nu(X_{0})}\otimes-d\overline{x}_{0}$
$=-\pi^{e\nu(X_{0})}\otimes\bigl(\overline{C}\,d\overline{w}\bigr)$
$=-C_{0}\otimes d\overline{w}\,=\,\alpha^{p}\lambda^{-p}\otimes d\overline{w}.$
This matches the formula in Proposition 8.3 2). ∎
Remark 8.5.
In Case 1, we cannot compute the Witt vector $(X_{0})$ associated with the Kummer class $u$, since
$X_{0}=\frac{1}{p\lambda}\ln(u),$
and $\ln(u)$ is not defined when $u\in\mathcal{O}_{\mathbb{K}}^{\times}$ with $\overline{u}\notin\kappa^{p}$. Nevertheless, we may differentiate formally, writing $d\!\ln(u)=\tfrac{du}{u}$. In the filtered module $\mathfrak{p}_{\mathbb{K}}^{-\frac{p}{p-1}}\otimes\Omega^{1}_{\kappa}$, we obtain
$dX_{0}\otimes 1=-\frac{1}{p\lambda}\frac{du}{u}\otimes 1.$
The relation (47) implies
$\frac{1}{p\lambda}\equiv-\lambda^{-p}\pmod{\mathfrak{p}_{\mathbb{K}}^{\frac{p+1}{p-1}}}$.
Substituting yields the reduction
$dX_{0}\otimes 1=\lambda^{-p}\otimes\frac{d\overline{u}}{\overline{u}},$
which agrees with Proposition 8.3 1).
Taken together, these computations indicate that the description provided by Theorem 8.1 should admit a mixed-characteristic analogue, with the Witt vector $(X_{0},\ldots,X_{s-1})$ playing the role of $(x_{0},\ldots,x_{s-1})$. Developing such an extension would offer a unified perspective on refined Swan conductors across characteristics. This direction is the subject of our ongoing work.